HUBBLE SPACE TELESCOPE DAILY REPORT #2244 PERIOD COVERED: 0000Z (UTC) 11/04/98 - 0000Z (UTC) 11/05/98 Daily Status Report as of 309/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As described in 2.1 and HSTAR 6757, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS, possibly affecting two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.3 Completed Two Sets of NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed Six Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.5 Completed WF/PC-2 7278 (The Connection Between The Obscuring Torus and the Masing Disk in H_20 Me) The WF/PC-2 was used to investigate the connection between the molecular torus and the emission line gas in a complete sample of H_2O megamasers. The proposal completed nominally with no reported problems. 1.6 Completed STIS/CCD/MA1/MA2 7673 (STIS MAMA Sensitivity and Focus Monitor) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to make observations of standard star GRW+70D582 in order to perform a MAMA sensitivity and focus monitoring program. This is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 7656 (Spectroscopic and Imaging Sensitivity) The Space Telescope Imaging Spectrograph (CCD) was used to make sensitivity measurements for all supported CCD imaging and spectroscopic modes. It is run once a year. The proposal executed as planned with no problems. 1.8 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed Seven Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. As described in 2.1 and HSTAR 6757, the acquisition for final iteration of this proposal defaulted to fine lock backup on one FGS, possibly affecting three observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.10 Completed NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD/MA1 7513 (The Stellar Content of Giant HII Regions) The Space Telescope Imaging Spectrograph (CCD/MA1) was used to obtain spectroscopy of two giant HII regions over the wavelength range 1200 Angstrom -- 10000 Angstrom. The proposal completed with no reported problems. 1.12 Completed Three Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed NIC/1/2/3 7827 (The Structure of Circumbinary Material in Multiple T-Tauri Systems) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class system GH-Tauri to study the structure of its circumbinary material. Recent studies of T-Tauri populations in nearby star forming clouds have shown that the majority of stars exist in multiple {binary or higher} systems. As described in 2.1 and HSTAR 6756, the acquisition for this proposal defaulted to fine lock backup on one FGS, possibly affecting seven observations. Otherwise, the observations were completed as planned, and no further problems were reported. 1.14 Completed WF/PC-2 6681 (Evolution of Pre-Main Sequence Circumstellar Nebulosity) The WF/PC-2 was used to make observations of DM-Tau to study the evolution of pre-main sequence circumstellar nebulosity. Initial HST results from imaging of disks around pre-main sequence stars provide compelling justification to observe a broader sample of young stars. The observations were completed as planned, and no anomalies were reported. 1.15 Completed Three Sets of WF/PC-2 6666 (The Mass Function and Binary Star Fraction of the Trapezium Cluster) The WF/PC-2 was used to take deep images of the Trapezium cluster to determine the low-mass end of the initial mass function, to search for brown dwarfs, and to explore the close binary population in the cluster. The observations completed as planned with no problems. 1.16 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 6673 (Black Holes and Gas Disks in a Complete Sample of Radio Loud UGC Ellipticals) The WF/PC-2 was used to make observations of radio load, elliptical galaxy UGC08419 and in order to search for possible black holes and gas disks. Hydrogen-Alpha and broad band observations were taken to determine the luminosity profiles and stellar populations (on the tens of parsec scale), and detect nuclear regions or disks of gas and dust. As described in 2.1 and HSTAR 6757, the acquisition for this proposal defaulted to fine lock backup on one FGS, possibly affecting seven observations. Otherwise, the observations were completed as planned, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 16 Successful : 16 The acquisition at 308/145000Z defaulted to fine lock backup on FGS-3 only when the scan step limit was exceeded on FGS-2. HSTAR 6756 was written. The proposal described in 1.13 may have been affected. As documented in HSTAR 6757, the acquisition at 308/223948Z defaulted to fine lock backup on FGS-3 due to FGS-1. The proposals described in 1.2, 1.9, and 1.17 may have been affected. Scheduled Reacquisitions : 2 Successful : 2 2.2 FHST Updates: Scheduled : 36 Successful : 36 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. The NSSC-1 status buffer was dumped at 308/2135Z per ROP NS-3. The Battery #2 capacity check test was initiated at 308/1327Z. Battery discharge began at 309/0005Z. ESTR-2 reconditioning was performed at 309/0914Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH